OpenStudy (anonymous):

Make "r" the subject of the formula V = (4/3) (pi) (r^3)

myininaya (myininaya):

\[V=\frac{4}{3}*\pi*r^3\] \[V*\frac{3}{4}*\frac{1}{\pi}=\frac{4}{3}*\pi*r^3*\frac{3}{4}*\frac{1}{\pi}\] \[\frac{3V}{4\pi}=r^3\] \[r^3=\frac{3V}{4\pi}\] \[(r^3)^\frac{1}{3}=(\frac{3V}{4\pi})^\frac{1}{3}\] \[r=(\frac{3V}{4\pi})^\frac{1}{3}\]

OpenStudy (anonymous):

Ur a beast !

myininaya (myininaya):

no u